OpenStudy (anonymous):

Which statement most accurately explains why x^6/x^7 is equal to x^3? A. two of the five x's on top cancel with the three x's on bottom, thus x^3 B. three of the five x's on top cancel with the three x's on bottom, thus x^3 C. two of the five x's on top cancel with the two x's on bottom, leaving three x's on top, thus x^3 D. three of the five x's on top cancel with the three x's on bottom, leaving three x's on top, thus x^3 help?

OpenStudy (anonymous):

at the risk of repeating myself, \[\frac{x^6}{x^7}\neq x^3\]

OpenStudy (anonymous):

I don't get what you mean mannnnnn >:(

OpenStudy (anonymous):

\[\frac{x^6}{x^7}=\frac{1}{x^{7-6}}=\frac{1}{x}\] so there is a mistake in the question

OpenStudy (anonymous):

lemme fix it